OpenStudy (anonymous):

Which of these solutions satisfies the Mean Value Theorem for the function the square root of (x) -1 in [1, 9]?

OpenStudy (anonymous):

\[f(x)=\sqrt{x}-1\]?

OpenStudy (anonymous):

\[f(9)=2\] \[f(1)=0\] \[\frac{2-0}{9-1}=\frac{2}{8}=\frac{1}{4}\]

OpenStudy (anonymous):

\[f'(x)=\frac{1}{2\sqrt{x}}\] set \[\frac{1}{2\sqrt{c}}=\frac{1}{4}\] and solve for c via \[2\sqrt{c}=4\] \[\sqrt{c}=2\] \[c=4\]

OpenStudy (anonymous):

yeahhh. thanks (:

OpenStudy (anonymous):

i don't know what "these solutions" are, but i hope 4 is one of them

OpenStudy (anonymous):

agree